OpenStudy (anonymous):

It has been quite some time since I've had to do this type of a calculation and I'm hoping that someone out there will know how to do this equation. 100 + 110/(1+r) = 1/(1+r) + 232/(1+r)2(squared)

hero (hero):

\[\frac{110}{1+r} = \frac{1}{1+r} + \frac{232}{(1+r)^2}\]

hero (hero):

Subtract 1/1+r from both sides: \[\frac{109}{1+r} = \frac{232}{(1+r)^2}\]

hero (hero):

Cross Multiply: 232(1+r) = 109(1+r)^2 Continue Solving from there...

OpenStudy (anonymous):

That's my problem, I haven't done this in over 20 years, and my brain just isn't computing it.

OpenStudy (anonymous):

then 109(1+r)=232\[109r=123\]\[r=123/109\]

hero (hero):

You divided by (1+r) which is illegal when isolating r

hero (hero):

You can't cancel variables when trying to isolate them, else you eliminate a solution

hero (hero):

I ended up with r^2 - 230r-122 = 0

hero (hero):

You can solve for r from there now.

OpenStudy (anonymous):

What about the first 100?